Calculate the molar solubility of CaF2 at 25 C in a solution that is 0.10 M in Ca(NO3)2.

Ksp = 3.9 E-11 for CaF2.

A. 3.9 E-7

B. 1.96 E-9

C. 7.8 E-10

D. 3.9 E-11
Answer #1

Here is the solution of your question. If you have any doubt or need any clarification please comment in comment box and will definitely resolve your query. If you find useful please upvote it. Thanks in advance.

Cafe & Ca 2+ + 26 11 S 25 Ca(NO3), = Ca2+ + 2003- 0.10.2 001 [Ca2+] = 0oots [F-] = 28 Ksp = [ca2+][F-32 Ksp = 0.10+5) (25) as